If theta is a third quadrant angle and cos theta= -2/3 , find sin theta.
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

cosine is adjacent over hypotenuse or 2/3
sine is sqrt (5) so that sqrt(5)^2+2^2=3^2
third quadrant, so sine is negative
sin theta= -sqrt (5)/3 ANSWER

Answer by ikleyn(52798)

.
I just answered this question under this link

https://www.algebra.com/algebra/homework/Trigonometry-basics/Trigonometry-basics.faq.question.1138392.html

https://www.algebra.com/algebra/homework/Trigonometry-basics/Trigonometry-basics.faq.question.1138392.html


Do not post duplicates . . . P L E A S E . . . .